OpenStudy (anonymous):

Part 1: Change 3x + 2y = 10 to slope-intercept form. You must show all work. Part 2: State the slope and y-intercept

OpenStudy (anonymous):

2y=10-3x y=(-3x+10)/2 y=(-3/2)x+5 Slope is -3/2 and y-intercept is 5

OpenStudy (anonymous):

All good?